OpenStudy (anonymous):

3. What is the value of -m if a) m = -1.4? b) m = 8?

OpenStudy (anonymous):

@DanJS @ShadowLegendX

OpenStudy (anonymous):

@ParthKohli @dan815

OpenStudy (anonymous):

we simply multiply both sides by -1 in each part

OpenStudy (anonymous):

\[m=-1.4\] multiply both sides by -1 \[m \times-1=-1.4\times-1\] \[-m=-\left( -1.4 \right)=+1.4\] thus \[-m=1.4\]

OpenStudy (anonymous):

: )

OpenStudy (anonymous):

so same with b it's -m = -8 right @chris00

OpenStudy (anonymous):